# Beamforming and Waveform Optimization for RF Wireless Power Transfer with Beyond Diagonal Reconfigurable Intelligent Surfaces

### v1

@@ Abstract
Radio frequency (RF) wireless power transfer (WPT) is a promising technology to seamlessly charge low-power devices, but its low end-to-end power transfer efficiency remains a critical challenge. To address the latter, low-cost transmit/radiating architectures, e.g., based on reconfigurable intelligent surfaces (RISs), have shown great potential. Beyond diagonal (BD) RIS is a novel branch of RIS offering enhanced performance over traditional diagonal RIS (D-RIS) in wireless communications, but its potential gains in RF-WPT remain unexplored. Motivated by this, we analyze a BD-RIS-assisted single-antenna RF-WPT system to charge a single rectifier, and formulate a joint beamforming and multi-carrier waveform optimization problem aiming to maximize the harvested power. We propose two solutions relying on semi-definite programming for fully connected BD-RIS and an efficient low-complexity iterative method relying on successive convex approximation. Numerical results show that the proposed algorithms converge to a local optimum and that adding transmit sub-carriers or RIS elements improves the harvesting performance. We show that the transmit power budget impacts the relative power allocation among different sub-carriers depending on the rectifier’s operating regime, while BD-RIS shapes the cascade channel differently for frequency-selective and flat scenarios. Finally, we verify by simulation that BD-RIS and D-RIS achieve the same performance under pure far-field line-of-sight conditions (in the absence of mutual coupling). Meanwhile, BD-RIS outperforms D-RIS as the non-line-of-sight components of the channel become dominant.
@@ I Introduction
Future wireless communication systems must ensure seamless green connectivity among numerous low-power devices. For this, it is essential to mitigate electronic waste resulting from battery replacements and reduce disruptions caused by battery depletion [1, 2, 3]. Energy harvesting (EH) technologies are fundamental enablers for this by providing wireless charging capability and promoting sustainable Internet of Things [4, 5]. EH devices may harvest energy from readily available or from dedicated sources. However, the former, a.k.a, ambient EH, may not be possible in all environments and/or may call for large device form factors [6], while the latter is supported by wireless power transfer (WPT) technologies such as inductive coupling, laser power beaming, magnetic resonance coupling, and radio frequency (RF) radiation. Hereinafter, we focus on RF-WPT, which can provide multi-user wireless charging capability over large distances while using the wireless communications infrastructure.
@@ I-A Preliminaries
A WPT system comprises three key building blocks: i) energy transmitter (ET); ii) wireless channel; and iii) energy receiver (ER). The charging signal is generated and amplified using a direct current (DC) power source at the ET. Then, it is upconverted to RF and transmitted over the wireless channel. The ER collects the signal and converts it to DC for EH. Each block includes some power consumption and loss sources, which bring new challenges to the system design. In fact, a key challenge of WPT systems is their inherently low end-to-end power transfer efficiency (PTE), which is given by
@@
Herein, $e_{1}$ is impacted by the power consumption sources at the transmitter side, such as high-power amplifier (HPA), while the channel losses, i.e., shadowing and fading, can decrease $e_{2}$. Note that the RF-to-DC conversion inefficiency at the receiver side influences $e_{3}$. Thus, it is important to address these and provide novel solutions to maximize efficiency.
@@
Waveform optimization is required to enhance efficiency at the ER. In fact, multitone waveforms can leverage the ER’s non-linearity and enhance the performance in terms of DC harvested power [7, 8]. Also, energy beamforming (EB) techniques can compensate for wireless channel losses by pointing the transmit signal effectively toward the ERs’ direction. The transmitter architecture determines the type of EB, which can be either active or passive. An active EB leverages active antenna elements connected to dedicated RF chains, while a passive EB utilizes low-cost nearly passive elements. For instance, a fully digital transmitter comprises active antenna elements, each with a dedicated RF chain, leading to high cost but offering the best beamforming gain. On the other hand, analog architectures relying, e.g., on phase shifters, can reduce the cost but also degrade the performance. There exist hybrid architectures, which combine the two mentioned architectures, leading to a reduced number of RF chains and a tradeoff between cost/complexity and performance [9].
@@
Traditional analog and hybrid architectures still need to cope with the challenges caused by complex analog networks. There are novel transmit architectures that avoid this additional complexity and provide EB gains at a much lower cost. For instance, dynamic metasurface antennas (DMA) utilize a limited number of RF chains, each one feeding multiple antenna elements on a waveguide [10, 11]. Moreover, reconfigurable intelligent surfaces (RIS) have become popular to enhance performance in wireless systems, especially in the presence of blockages/obstacles. This emerging technology may rely on planar surfaces comprising nearly passive scattering elements that introduce amplitude and phase changes to incident electromagnetic waves. Therefore, reflective-type RIS can smartly tune the reflected signal and point it toward the desired direction, providing significant EB gains [12]. In addition to reflective-type RIS, there exist some novel RIS-based architectures such as active RIS providing extra capabilities at the transmitter [13], or reflective-refractive models [14].
@@
A recent generalization of diagonal RIS (D-RIS) is given by beyond diagonal RIS (BD-RIS), which is characterized by scattering matrices not constrained to be diagonal, thus elements/ports interconnected via tunable impedances [15]. The optimality of BD-RIS in lossless designs was proved in [16]. BD-RIS, through reconfigurable inter-connections, enables impinging waves to flow through the surfaces, hence a new degree of freedom and higher flexibility to manipulate waves by forming fully-, group-, tree-, and forest-connected structures [16, 17]. The former provides the highest flexibility in shaping the reflected signal, though lower complexities based on tree- and forest-connected architectures are attractive to achieve the performance-complexity Pareto frontier [17, 18]. Note that decreasing the number of connections between the scattering elements reduces both the performance gains and complexity, showcasing an interesting performance-complexity trade-off. BD-RIS is a general framework enabling multiple modes, such as reflective, hybrid, transmissive and reflective, and multi-sector, while encompassing other RIS architectures as special cases [19, 20].
@@ I-B Prior Works
Waveform and EB optimization for WPT with non-linear EH has received significant research attention, although most works focus on traditional fully digital transmitters. For instance, the authors of [21] propose transmit and receive EB solutions for a fully digital multiple-input multiple-output (MIMO) system to increase the DC harvested power, demonstrating the superiority of RF over DC combining. The authors of [22] investigate the design of WPT waveforms for different EH models and show that waveforms designed considering nonlinear EH yield substantial improvements in harvested DC power. Interestingly, in [23], the authors propose a low-complexity channel-adaptive waveform design for single-ER WPT systems leading to near-optimal EH performance. In [24], a low-complexity near-optimal EB is proposed for maximizing the weighted sum harvested DC power in multi-ER scenarios. Moreover, the waveform and EB optimization problem for large-scale multi-antenna WPT systems has been investigated in [25], demonstrating the effectiveness of large-scale WPT deployments in improving the PTE. The authors of [26] leverage EB and optimized multi-sine waveforms in a MIMO WPT system and show that the joint optimization leads to significant gains compared to EB-only designs in terms of harvested power. All mentioned studies considered just the EH non-linearity, while [27] considers both HPA and rectifier’s non-linearity. Therein, it is shown that the waveform design considering HPA non-linearity leads to significant gains compared to the ones focusing only on the rectifier. Moreover, multiple works, e.g., [28, 29, 30], focus on RF power maximization by considering linear EH for fully digital WPT.
@@
Novel low-cost WPT transmit architectures for reducing the implementation cost/complexity is another topic gaining attention recently. The authors of [31] study the waveform and EB optimization for DMA-assisted WPT systems and show that DMA can be used to meet the DC requirement of the ERs with considerably fewer RF chains compared to a fully-digital counterpart. The authors of [32] propose a minimum-power beamforming design for simultaneous wireless information and power transfer (SWIPT) systems with a hybrid transmit architecture. The authors of [33] derive the number of RIS elements in an RIS-aided WPT system required to outperform traditional WPT relying on active antenna elements. Interestingly, the harvested power maximization problem is addressed in [34, 35] for RIS-aided WPT and SWIPT systems with non-linear EH, respectively. Therein, the authors highlight the extra beamforming gains on the harvested DC power provided by RISs, although only focusing on conventional D-RIS. Note that BD-RIS architectures have been widely investigated for communication purposes [36, 17, 16, 15, 37], while current research on RIS-aided WPT is still limited to conventional diagonal (local) RIS [35, 34]. Notably, another important yet unexplored topic is the use of BD-RIS for SWIPT systems with non-linear EH models. However, this work focuses on WPT alone, noting that maximizing the harvested DC power is first crucial as it directly impacts the achievable information transmission rate in SWIPT systems.
@@ I-C Contributions
Efficient WPT system design and optimization requires accurate non-linear EH models [8, 22]. However, most existing works employing non-linear EH primarily focus on traditional fully digital WPT designs. In contrast, this paper investigates BD-RIS-assisted WPT under non-linear EH models, a novel direction not yet explored. [^1] This work is an extension of our conference paper [38], wherein we present a BD-RIS EB and waveform optimization algorithm. In the current work, we propose two other solutions for BD-RIS beamforming and an efficient iterative solution for waveform optimization. Moreover, we provide thorough algorithm performance comparisons and a comprehensive waveform behavior analysis. Our main contributions are:
@@
First, we formulate a joint EB and waveform optimization problem for a single-input single-output (SISO) multi-carrier BD-RIS-aided WPT system to maximize the DC harvested power. For this, we assume there is no direct path between the ET and ER. Our problem requires novel solutions because i) the methods available in the literature for D-RIS [34] do not apply to BD-RIS due to the additional mathematical complexity caused by the non-diagonal scattering matrix; ii) our considered problem is different from existing ones in the BD-RIS literature, e.g., [36, 17], which focus on RF power maximization. The reason is that the DC harvested power depends on the fourth moment of the signal in addition to the second moment, a.k.a., RF power. Thus, the multi-carrier signal must be optimized such that the EH non-linearity is leveraged efficiently and the DC power is maximized [22].
@@
Second, we decouple the waveform optimization and EB problem via alternating optimization, and propose:
@@
i) An efficient iterative method relying on successive convex approximation (SCA) and Karush–Kuhn–Tucker (KKT) conditions for the waveform optimization problem.
@@
ii) An EB optimization framework by relying on SCA and semi-definite programming (SDP), which models the rank-1 constraint using the atomic and Frobenius norms [39] and guarantees solution feasibility.
iii) A semi-definite relaxation (SDR)-based EB method with lower time complexity compared to the one relying on SDP due to fewer approximations.
@@
iv) An efficient iterative EB method by leveraging the Neumann series approximation, SCA, and KKT conditions. Specifically, we approximate the problem using local convex subproblems, which can be solved using SCA. Then, we leverage the KKT conditions to solve each subproblem and an iterative method using low-complexity closed-form expressions to update the optimization variables. The computational complexity of the iterative method is reduced compared to the SDR-based approach from the fourth to the second degree of the number of RIS elements.
@@
i) The proposed algorithms converge to a local optimum solution, while the convergence time scales with the number of sub-carriers and RIS elements.
@@
iv) The frequency selectivity of the channel impacts the way that BD-RIS shapes the cascade channel, leading to performance changes. Specifically, when the channel tends to be frequency-flat, the BD-RIS can shape the cascade channel more effectively due to its frequency-flat nature. On the other hand, the variation among different sub-carriers of the cascade channel is higher for a frequency-selective case since BD-RIS impacts the channels at different sub-carriers differently.
@@
vi) BD-RIS does not introduce any additional performance gains in terms of harvested DC power compared to D-RIS in far-field line-of-sight (LoS) channels (in the absence of mutual coupling) even with multi-carrier signals. On the other hand, in Rician channels with non-LoS (NLoS) components, BD-RIS outperforms D-RIS by effectively tuning the cascade channel, leading to increased harvested DC power. This extends the findings of [16] to multi-carrier WPT.
@@
[…]

### v2

@@ Abstract
Radio frequency (RF) wireless power transfer (WPT) is a promising technology to seamlessly charge low-power devices, but its low end-to-end power transfer efficiency remains a critical challenge. To address the latter, low-cost transmit/radiating architectures, e.g., based on reconfigurable intelligent surfaces (RISs), have shown great potential. Beyond diagonal (BD) RIS is a novel branch of RIS offering enhanced performance over traditional diagonal RIS (D-RIS) in wireless communications, but its potential gains in RF-WPT remain unexplored. Motivated by this, we analyze a BD-RIS-assisted single-antenna RF-WPT system to charge a single rectifier, and formulate a joint beamforming and multi-carrier waveform optimization problem aiming to maximize the harvested power. We propose two solutions relying on semi-definite programming for fully connected BD-RIS, a successive convex approximation (SCA)-based beamforming approach, and an efficient low-complexity iterative method relying on SCA. Numerical results show that the proposed algorithms converge and that adding transmit sub-carriers or RIS elements improves the harvesting performance. We show that the transmit power budget impacts the relative power allocation among different sub-carriers depending on the rectifier’s operating regime, while BD-RIS shapes the cascade channel differently for frequency-selective and flat scenarios. Finally, we verify by simulation that BD-RIS and D-RIS achieve the same performance under pure far-field line-of-sight conditions (in the absence of mutual coupling). Meanwhile, BD-RIS outperforms D-RIS as the non-line-of-sight components of the channel become dominant.
@@ I Introduction
Future wireless communication systems must ensure seamless green connectivity among numerous low-power devices. For this, it is essential to mitigate electronic waste resulting from battery replacements and reduce disruptions caused by battery depletion [intro1, intro2, lópez2023highpower]. Energy harvesting (EH) technologies are fundamental enablers for this by providing wireless charging capability and promoting sustainable Internet of Things [3gppamb, ZEDHEXA]. EH devices may harvest energy from readily available or dedicated sources. However, the former, a.k.a., ambient EH, may not be possible in all environments and/or may call for large device form factors [intro3], while the latter is supported by wireless power transfer (WPT) technologies such as inductive coupling, laser power beaming, magnetic resonance coupling, and radio frequency (RF) radiation. Hereinafter, we focus on RF-WPT, which can provide multi-user wireless charging capability over large distances while using the wireless communications infrastructure.
@@ I-A Preliminaries
A WPT system comprises three key building blocks: i) energy transmitter (ET); ii) wireless channel; and iii) energy receiver (ER). The charging signal is generated and amplified using a direct current (DC) power source at the ET. Then, it is upconverted to RF and transmitted over the wireless channel. The ER collects the signal and converts it to DC for EH. Each block includes some power consumption and loss sources, which bring new challenges to the system design. In fact, a key challenge of WPT systems is their inherently low end-to-end power transfer efficiency, which is given by
@@
Herein, $e_{1}$ is impacted by the power consumption sources at the transmitter side, such as high-power amplifier, while the channel losses, i.e., shadowing and fading, can decrease $e_{2}$. Note that the RF-to-DC conversion inefficiency at the receiver side influences $e_{3}$. Thus, it is important to address these and provide novel solutions to maximize efficiency.
@@
Waveform optimization is required to enhance efficiency at the ER. In fact, multi-tone waveforms can leverage the ER’s non-linearity and enhance the performance in terms of DC harvested power [clerckx-nonlinearharvester, clerckx2018beneficial]. Also, beamforming techniques can compensate for wireless channel losses by pointing the transmit signal effectively toward the ERs’ direction. The transmitter architecture determines the type of beamforming, which can be either active or passive. An active beamforming leverages active antenna elements connected to dedicated RF chains, while a passive beamforming utilizes low-cost nearly passive elements. For instance, a fully digital transmitter comprises active antenna elements, each with a dedicated RF chain, leading to high cost but offering the best beamforming gain. On the other hand, analog architectures relying, e.g., on phase shifters, can reduce the cost but also degrade the performance. There exist hybrid architectures, which combine the two mentioned architectures, leading to a reduced number of RF chains and a tradeoff between cost/complexity and performance [hybridbeamsurvey].
@@
Traditional analog and hybrid architectures still need to cope with the challenges caused by complex analog networks. There are novel transmit architectures that avoid this additional complexity and provide beamforming gains at a much lower cost. For instance, dynamic metasurface antennas (DMA) utilize a limited number of RF chains, each one feeding multiple antenna elements on a waveguide [DMAWPT, MyEBDMA]. Moreover, reconfigurable intelligent surfaces (RIS) have become popular to enhance performance in wireless systems, especially in the presence of blockages/obstacles. This emerging technology may rely on planar surfaces comprising nearly passive scattering elements that introduce amplitude and phase changes to incident electromagnetic waves. Therefore, reflective-type RIS can smartly tune the reflected signal and point it toward the desired direction, providing significant beamforming gains [RIS-basis]. In addition to reflective-type RIS, there exist some novel RIS-based architectures such as active RIS providing extra capabilities at the transmitter [active_ris], or reflective-refractive models [reflect_refract_RIS].
@@
A recent generalization of diagonal RIS (D-RIS) is given by beyond diagonal RIS (BD-RIS), which is characterized by scattering matrices not constrained to be diagonal, thus elements/ports interconnected via tunable impedances [bdris_main_ref]. The optimality of BD-RIS in lossless designs was proved in [BD-RIS_scattering_clerckx]. BD-RIS, through reconfigurable inter-connections, enables impinging waves to flow through the surfaces, hence a new degree of control and higher flexibility to manipulate waves by forming fully-, group-, tree-, and forest-connected structures [BD-RIS_scattering_clerckx, BD-RIS_General]. The former provides the highest flexibility in shaping the reflected signal, though lower complexities based on tree- and forest-connected architectures are attractive to achieve the performance-complexity Pareto frontier [BD-RIS_General, Pareto_Frontier_nerini]. Note that decreasing the number of connections between the scattering elements reduces both the performance gains and complexity, showcasing an interesting performance-complexity trade-off. BD-RIS is a general framework enabling multiple modes, such as reflective, hybrid, transmissive and reflective, and multi-sector, while encompassing other RIS architectures as special cases [Multi_Sector_BDRIS, TR_REF_BDRIS].
@@ I-B Prior Works
Waveform and beamforming optimization for WPT with non-linear EH has received significant research attention, although most works focus on traditional fully digital transmitters. For instance, the authors of [BFRFDCsingletone] propose transmit and receive beamforming solutions for a fully digital multiple-input multiple-output (MIMO) system to increase the DC harvested power, demonstrating the superiority of RF over DC combining. The authors of [clreckxWFdesign] investigate the design of WPT waveforms for different EH models and show that waveforms designed considering nonlinear EH yield substantial improvements in harvested DC power. Interestingly, in [brunolowcomp], the authors propose a low-complexity channel-adaptive waveform design for single-ER WPT systems, leading to near-optimal EH performance. In [azarbahram2024deepreinforcementlearningmultiuser], a low-complexity near-optimal beamforming is proposed for maximizing the weighted sum harvested DC power in multi-ER scenarios. Moreover, the waveform and beamforming optimization problem for large-scale multi-antenna WPT systems has been investigated in [largescale], demonstrating the effectiveness of large-scale WPT deployments in improving the power transfer efficiency. The authors of [jointWFandBFMIMOclreckx] leverage beamforming and optimized multi-sine waveforms in a MIMO WPT system and show that the joint optimization leads to significant gains compared to beamforming-only designs in terms of harvested power. All mentioned studies considered just the EH non-linearity, while [SISOAllClreckx] considers both the power amplifier and rectifier’s non-linearity. Therein, it is shown that the waveform design considering power amplifier non-linearity leads to significant gains compared to those focusing only on the rectifier. Moreover, multiple works, e.g., [OnelRadioStripes, onellowcomp, azarbahram2023radio], focus on RF power maximization by considering linear EH for fully digital WPT.
@@
Novel low-cost WPT transmit architectures for reducing the implementation cost/complexity is another topic gaining attention recently. The authors of [azarbahram2024waveform] study the waveform and beamforming optimization for DMA-assisted WPT systems and show that DMA can be used to meet the DC requirement of the ERs with considerably fewer RF chains compared to a fully-digital counterpart. The authors of [hybrid_SWIPT] propose a minimum-power beamforming design for simultaneous wireless information and power transfer (SWIPT) systems with a hybrid transmit architecture. The authors of [ris_rf_power] derive the number of RIS elements in an RIS-aided WPT system required to outperform traditional WPT relying on active antenna elements. Interestingly, the harvested power maximization problem is addressed in [clerckxRISWPT, ris_swipt_nonlin] for RIS-aided WPT and SWIPT systems with non-linear EH, respectively. Therein, the authors highlight the extra beamforming gains on the harvested DC power provided by RISs, although only focusing on conventional D-RIS. Note that BD-RIS architectures have been widely investigated for communication purposes [BD-RIS_Graph, BD-RIS_General, BD-RIS_scattering_clerckx, bdris_main_ref, direnzoclerckxmutual], while current research on RIS-aided WPT is still limited to conventional diagonal (local) RIS [ris_swipt_nonlin, clerckxRISWPT]. Notably, another important yet unexplored topic is the use of BD-RIS for SWIPT systems with non-linear EH models. However, this work focuses on WPT alone, noting that maximizing the harvested DC power is first crucial as it directly impacts the achievable information transmission rate in SWIPT systems.
@@ I-C Contributions
Efficient WPT system design and optimization requires accurate non-linear EH models [clerckx2018beneficial, clreckxWFdesign]. However, most existing works employing non-linear EH primarily focus on traditional fully digital WPT designs. In contrast, this paper investigates BD-RIS-assisted WPT under non-linear EH models, a novel direction not yet explored. [^1] This work is an extension of our conference paper [myBDrisWCNC], wherein we present a BD-RIS beamforming and waveform optimization algorithm. In the current work, we propose two other solutions for BD-RIS beamforming and an efficient iterative solution for waveform optimization. Moreover, we provide thorough algorithm performance comparisons and a comprehensive waveform behavior analysis. Our main contributions are:
@@
First, we formulate a joint beamforming and waveform optimization problem for a single-input single-output (SISO) multi-carrier BD-RIS-aided WPT system to maximize the DC harvested power. For this, we assume there is no direct path between the ET and ER. Our problem requires novel solutions due to the following reasons:
i) The methods available in the literature for WPT with D-RIS [clerckxRISWPT] do not apply to BD-RIS due to the additional mathematical complexity caused by the non-diagonal scattering matrix. Specifically, when optimizing BD-RIS, in contrast to D-RIS, all the elements of the scattering matrix can be non-zero, and a unitary constraint must be forced on the designed matrix, which significantly increases the mathematical complexity and calls for novel mathematical solutions.
ii) The methods available for optimizing BD-RIS, e.g., as in [BD-RIS_Graph, BD-RIS_General], which focus on RF power maximization, do not apply to the WPT case with non-linear EH. The reason is that the DC harvested power depends on the fourth moment of the signal in addition to the second moment, a.k.a., RF power. Specifically, RF power maximization only deals with the complexity caused by coupling between the waveform and beamforming. On the other hand, DC power maximization introduces inter-subcarrier dependencies in multi-carrier waveform and beamforming design, requiring novel methods to exploit EH non-linearity effectively [clreckxWFdesign].
@@
Second, we decouple the waveform optimization and beamforming problem via alternating optimization, and propose:
@@
i) An efficient iterative method relying on successive convex approximation (SCA) and Karush–Kuhn–Tucker (KKT) conditions for the waveform optimization problem. Specifically, we prove that the objective function is convex w.r.t the waveform, hence we approximate it using its first-order Taylor approximation. The resulting problem is convex in the local neighborhood, and a closed-form solution can be obtained using KKT conditions and can then be iteratively updated.
@@
ii) A beamforming optimization framework relying on SCA and semi-definite programming (SDP), which models the rank-1 constraint using the atomic and Frobenius norms [shabir2024electromagneticallyconsistentoptimizationalgorithms]. Unlike semi-definite relaxation (SDR), this approach guarantees a rank-1 feasible solution without relaxation.
iii) SDR-based beamforming method, which relaxes the rank-1 constraint, but has lower time complexity compared to the one relying on SDP due to fewer approximations. The proposed SDP and SDR-based methods differ fundamentally from [clerckxRISWPT], as they address the added complexities introduced by the non-diagonal structure of the BD-RIS scattering matrix and the imposed unitary constraint. Notably, the SDP-based method further distinguishes itself by preserving the rank-1 constraint, in contrast to SDR-based approaches that relax this constraint.
iv) SCA-based beamforming, which leverages Taylor expansion to approximate the problem by dividing the unitary constraint into two separate regions. This algorithm has lower computational complexity than SDR, but yields a suboptimal solution due to the constraint relaxation.
@@
v) An efficient iterative beamforming method by leveraging the Neumann series approximation, SCA, and KKT conditions. Specifically, we approximate the problem with local convex subproblems by leveraging the Neumann series. Then, we leverage the KKT conditions to solve each subproblem and use an iterative method using low-complexity closed-form expressions to update the optimization variables. The final solution is iteratively obtained using the SCA method. This method reduces the computational complexity from the fourth to the second degree of the number of RIS elements compared to the SDP and SDR-based approaches.
@@
i) The proposed algorithms converge, with convergence time scaling with the number of sub-carriers and RIS elements.
@@
iv) The frequency selectivity of the channel impacts how BD-RIS shapes the cascade channel, leading to performance changes. Specifically, when the incident and reflective channels tend to be frequency-flat, the BD-RIS can shape the cascade channel more effectively since it affects all the sub-carriers similarly. On the other hand, in the frequency-selective case, the variation among sub-carriers in the cascade channel is greater because the BD-RIS configuration affects each sub-carrier differently.
@@
vi) BD-RIS does not introduce any additional performance gains in terms of harvested DC power compared to D-RIS in far-field line-of-sight (LoS) channels (in the absence of mutual coupling) even with multi-carrier signals. On the other hand, in Rician channels with non-LoS (NLoS) components, BD-RIS outperforms D-RIS by effectively tuning the cascade channel, leading to increased harvested DC power. This extends the findings of [BD-RIS_scattering_clerckx] to multi-carrier WPT.
@@
Notations: Bold lower-case and upper-case letters represent vectors and matrices. The $\ell_{2}$-norm operator is denoted by $\left\lVert\cdot\right\rVert$, while $\left\lVert\cdot\right\rVert_{\star}$ and $\left\lVert\cdot\right\rVert_{F}$ are the atomic and Frobenius norm, respectively. $\Re\{\cdot\}$ and $\Im\{\cdot\}$ denote the real and imaginary parts of a complex matrix or scalar, and $|\cdot|$ represents the absolute value operator. Moreover, $(\cdot)^{T}$, $(\cdot)^{H}$, and $(\cdot)^{\star}$ denote the transpose, transposed conjugate, and conjugate, respectively. Additionally, $\cdot]_{i,l}$ denotes the element in the $l$th column and the $i$th row of a matrix. The vectorization operator is represented by $\mathrm{Vec}(\cdot)$, and its inverse is denoted by $\mathrm{Vec}^{-1}_{D\times D}(\cdot)$}. Finally, $\mathbf{I}_{D}$ represents a $D\times D$ identity matrix and $\mathrm{diag}(\mathbf{a})$ is a diagonal matrix with its main diagonal being the elements of vector $\mathbf{a}$.